OpenStudy (anonymous):

Tough one! Solve the equation for x. a2x2 + 14ax + 49 = 0 (a ≠ 0) What is the: x = _____ (smaller value) x = _____ (larger value)

OpenStudy (anonymous):

a= +/- 7

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

if by a2x2 you mean a^2 yes.

OpenStudy (anonymous):

x=-7/a only one solution

OpenStudy (anonymous):

oh woops i see the problem ok hang on... (ax+7)^2=0 so -7/a is right.

myininaya (myininaya):

(ax+7)^2=0 ax+7=0 ax=-7 x=-7/a a does not equal 0

myininaya (myininaya):

as a gets closer to 0 from the right, you are geting closer to the smallest possible value for a but you will never reach the smallest value since -7/a has vertical asymptote at a=0 now the greates possible value will also never be reach because of a=0